# Recursive numeral systems are highly regular and easy to process

### v1

@@ 7 Limitations
The next limitation also concerns the data. D&S used the language sample from the World’s Atlas of Language Structures (WALS) (Comrie et al., 2013). While, as they pointed out, WALS was constructed with a focus on language diversity, we found that the sample comprises many languages from large language families such as Indo-European and Austronesian. This could be taken as “representative” in so far as the sample is roughly proportional to the number of languages across families, but a more appropriate sample might be one with more balance across large and small families, and is controlled for geographical effects (Dryer, 1989; Ladd et al., 2015).
@@
Appendix A Genetic Algorithm of Denić and Szymanik (2024)
In this Appendix we describe the Genetic Algorithm used by Denić and Szymanik (2024) to generate the Pareto frontier of optimal recursive numeral systems (Algorithm 1). Note that D&S to evaluate a $(D,M)$ pair utilise Hurford’s grammar to generate numerals for each number from 1991-99, keeping the shortest numeral for each of them. Yang and Regier (2025) utilise a similar algorithm but includes also suppletives $S$ beside digits and multipliers, use the constrains over Hurford’s grammar discussed in Section 2. While D&S perform mutations only on the sets $D$ and $M$, Y&R includes the suppletives $S$ and exception rules.

### v2

@@ 7 Limitations
The next limitation also concerns the data. D&S used the language sample from the World’s Atlas of Language Structures (WALS) (wals-s1). While, as they pointed out, WALS was constructed with a focus on language diversity, we found that the sample comprises many languages from large language families such as Indo-European and Austronesian. This could be taken as “representative” in so far as the sample is roughly proportional to the number of languages across families, but a more appropriate sample might be one with more balance across large and small families, and is controlled for geographical effects (dryer1989large; laddCorrelationalStudiesTypological2015).
@@
Appendix A Genetic Algorithm of denic2024recursive
In this Appendix we describe the Genetic Algorithm used by denic2024recursive to generate the Pareto frontier of optimal recursive numeral systems (Algorithm 1). Note that D&S to evaluate a $(D,M)$ pair utilise Hurford’s grammar to generate numerals for each number from 1991-99, keeping the shortest numeral for each of them. yang2025re utilise a similar algorithm but includes also suppletives $S$ beside digits and multipliers, use the constrains over Hurford’s grammar discussed in Section 2. While D&S perform mutations only on the sets $D$ and $M$, Y&R includes the suppletives $S$ and exception rules.